# TSKANMixer: Kolmogorov–Arnold Networks with MLP-Mixer Model for Time Series Forecasting

### v1

@@ Introduction
In this paper, we explore and evaluate the application of a KAN layer to the MLP-based TSMixer architecture. We introduce three architectures of TSKANMixer as illustrated in Figure 2. The proposed models apply the KAN framework to learn complex, non-linear relationships in temporal data. The first proposed architecture, presented in Figure 2(a), uses KAN for temporal projection on the time domain as an alternative to a fully-connected layer in TSMixer (Chen et al. 2023). It maps the time series from the input length $L$ to the forecast horizon $H$ by learning the complex relationships between past inputs and future predictions. The second proposed architecture, presented in Figure 2(b), extends TSMixer by adding a new KAN-based time mixing layer between mixer layers and temporal projection to intensify the capability to uncover the temporal patterns in time series. All architectures use a two-depth KAN layer.
@@
TSKANMixer (v01) TSKANMixer (v02) TSMixer dataset L H time/epoch (sec) training time (sec) time/epoch (sec) training time (sec) timee/poch (sec) training time (sec) ETTh1 512 96 21.29 4885.08 63.70 11869.71 4.10 263.01 ETTh2 512 96 40.69 12276.67 63.59 19302.69 4.74 312.52 ETTm1 512 96 88.81 11662.25 264.63 39819.87 26.83 1282.46 ETTm2 512 96 171.43 19242.63 263.53 63953.10 27.26 2032.07 NN5_daily 56 56 25.82 1884.76 22.80 1869.92 0.62 135.86 NN5_weekly 16 8 6.19 1354.61 3.21 125.01 0.08 28.89 CIF_2016 24 12 2.97 204.98 4.03 454.99 0.07 59.96 Hospital 24 12 20.32 589.23 15.09 2489.44 0.04 11.9 Exchange 60 30 5.04 181.33 9.70 339.55 10.18 2421.82 FRED_MD 48 12 30.89 1730.17 22.46 1639.77 0.33 181.19
@@
TSKANMixer (v02) dataset L H Batch Blocks Dropout Hidden size Learing rate KAN_dim KAN_grid KAN_k ETTh1 512 96 320 2 0.3 64 0.0001 1025 5 3 ETTh2 512 96 320 2 0.3 64 0.0001 1025 5 3 ETTm1 512 96 320 2 0.3 64 0.0001 1025 5 3 ETTm2 512 96 320 2 0.3 64 0.0001 1025 5 3 NN5_daily 56 56 16 4 0.9 32 0.001 14 2 3 NN5_weekly 16 8 8 6 0.7 32 0.001 8 7 3 CIF_2016 24 12 16 4 0.4 24 0.001 12 7 3 Hospital 24 12 8 6 0.3 16 0.001 3 10 2 Exchange 60 30 32 2 0.3 16 0.001 15 10 2 FRED_MD 48 12 16 2 0.5 16 0.001 0 5 2

### v2

@@ Introduction
In this paper, we explore and evaluate the application of a KAN layer to the MLP-based TSMixer architecture. We introduce two architectures of TSKANMixer as illustrated in Figure 2. The proposed models apply the KAN framework to learn complex, non-linear relationships in temporal data. The first proposed architecture, presented in Figure 2(a), uses KAN for temporal projection on the time domain as an alternative to a fully-connected layer in TSMixer (Chen et al. 2023). It maps the time series from the input length $L$ to the forecast horizon $H$ by learning the complex relationships between past inputs and future predictions. The second proposed architecture, presented in Figure 2(b), extends TSMixer by adding a new KAN-based time mixing layer between mixer layers and temporal projection to intensify the capability to uncover the temporal patterns in time series. All architectures use a two-depth KAN layer.
@@
TSKANMixer (v01) TSKANMixer (v02) TSMixer dataset L H time/epoch (sec) training time (sec) time/epoch (sec) training time (sec) timee/poch (sec) training time (sec) ETTh1 512 96 21.29 4885.08 63.70 11869.71 4.10 263.01 ETTh2 512 96 40.69 12276.67 63.59 19302.69 4.74 312.52 ETTm1 512 96 88.81 11662.25 264.63 39819.87 26.83 1282.46 ETTm2 512 96 171.43 19242.63 263.53 63953.10 27.26 2032.07 NN5_daily 56 56 25.82 1884.76 22.80 1869.92 0.62 135.86 NN5_weekly 16 8 6.19 1354.61 3.21 125.01 0.08 28.89 CIF_2016 24 12 2.97 204.98 4.03 454.99 0.07 59.96 Hospital 24 12 20.32 589.23 15.09 2489.44 0.04 11.9 Exchange 60 30 5.04 181.33 9.70 339.55 10.18 2421.82 FRED_MD 48 12 30.89 1730.17 19.98 3036.34 0.33 181.19
@@
TSKANMixer (v02) dataset L H Batch Blocks Dropout Hidden size Learing rate KAN_dim KAN_grid KAN_k ETTh1 512 96 320 2 0.3 64 0.0001 1025 5 3 ETTh2 512 96 320 2 0.3 64 0.0001 1025 5 3 ETTm1 512 96 320 2 0.3 64 0.0001 1025 5 3 ETTm2 512 96 320 2 0.3 64 0.0001 1025 5 3 NN5_daily 56 56 16 4 0.9 32 0.001 14 2 3 NN5_weekly 16 8 8 6 0.7 32 0.001 8 7 3 CIF_2016 24 12 16 4 0.4 24 0.001 12 7 3 Hospital 24 12 8 6 0.3 16 0.001 3 10 2 Exchange 60 30 32 2 0.3 16 0.001 15 10 2 FRED_MD 48 12 16 2 0.5 16 0.001 5 5 2